suppose the roots of ax^2 + bx + c =0 are r and s. which one of the following has roots ar + b and as + b ?
(A) x^2 - bx - ac =0
(B) x^2 - bx + ac =0
(C) x^2 + 3bx + ca + 2b^2 =0
(D) x^2 + 3bx - ca + 2b^2 =0
(E) not given

Answer by richard1234(7193)

Recall that, if z is a root of a polynomial, then x-z is a factor of the polynomial.

Hence, we can write our new polynomial equation as (x - (ar + b))(x - (as + b)) = 0. Expanding the LHS, this yields





Recall that by Vieta's formulas, r+s = -b/a and rs = c/a. We can substitute these expressions in:





b^2 cancels, so you will see that this equation is the same as in answer choice B.